Mathematics 122 – Fall 96

How to prepare for the Final Exam

Here are ten simple suggestion to keep in mind when preparing for the final.

1. Study one section of the material for the exam at a time. Make sure you understand the material in a section before you move to the next one.

2. Read the material from each section from the book and/or your class notes. Follow the order indicated in the summary of main topics.

3. Read the examples I did in class for each section.

4. Make for yourself a brief summary of the main concepts and/or formulas for that particular section (like the one I did in class before the first exam).

5. Re-do the homework problems assigned in that section and try a few extra ones from the book.

6. Do the quiz (if any) for that section without looking at your answers or the ones I gave you. Then check them.

7. After you did all of the above with all the sections, do again the in-class exams.

8. Do the practice problems I gave you and the practice exam sold at the bookstore.

9. Compare the kind of problems and questions asked in the practice problems and review the summary you made to make sure you did not left out any important concept.

10. Study hard. Put your last effort in the course and it will show in your grade. Don’t panic and, if possible, rest well before the exam.

The final exam will concentrate on the material not covered but the midterm, but you should know the material covered in the midterm very well by now and some questions about that material could be in the final.

The exam is on Wed., December 18, 7:00 - 10:00 pm Exam room: 2022 Learned

Make sure the batteries in your calculator are charged!!! Good luck.